OpenStudy (anonymous):

If f(x)= x/2 - 2 and g(x) = 2x^2 + x-3, find (f+g) (x) A. 2x^2- x/2 + 1 B. 2x^2+ 3/2x +1 C. x^2 - 6 D. 2x^2 +3/2x - 5

OpenStudy (anonymous):

Thank you for the help guys

OpenStudy (trojanpoem):

formulas : (g+f)(x) = f(x) + g(x) (g-f)(x) = g(x) - f(x) (f/g)(x) = f(x)/g(x) (f*g)(x) = f(x) * g(x) Regarding the problem , (f+g)(x) = f(x) + g(x) = 0.5x - 2 + 2x^2 +x -3 = 2x ^2 + 1.5 x - 5 (B)